Operations Report – Nov 8th

Crew 214 Operations Report 8-11-2019

SOL: 12

Name of person filing report: Andrew Wheeler

Non-nominal systems: Opportunity, Suit 10

Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning

Generator: run

Hours run: 13h 45m

From what time last night: 1745

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 76%

Diesel Reading – 55%

Station Propane Reading – 77%

Ethanol Free Gasoline: N/A

Water (loft tank): 20 gallons

Water Meter: 1456346 units

Water (static tank): 332 gallons

Static to Loft Pump used – no

Water in Green Hab: 225 gallons

Water in Science Dome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: not at hab

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Spirit rover used: no

Hours: 108.8

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 100%

Currently charging: yes

Opportunity rover used: no

Hours: 65.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: no

Curiosity rover used: no

Hours: 114.7

Beginning charge: (Before EVA) : 100%

Ending charge: (On return from EVA, before charging): 100%

Currently charging: yes

Notes on rovers: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no

Reason for use: N/A

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were used today: 0

Notes on ATVs: Assigned to the director

HabCar used and why, where? HabCar returned from service in Bicknell.

CrewCar used and why, where? CrewCar returned from service in Bicknell.

General notes and comments: N/A

Summary of internet: Functional

Summary of suits and radios: Nominal

Summary of Hab operations: All systems nominal

Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered. Additional seeds were catalogued in preparation for planting. GreenHab heater turned on automatically at 1700.

Summary of Science Dome operations: All systems nominal. Sample processing and experiments completed. All equipment used has been cleaned and returned to storage. Unused samples for ongoing experiments by Crew 215 placed in storage.

Summary of RAM operations: Not used

Summary of any observatory issues: Nominal

Summary of health and safety issues: All crew healthy

Questions, concerns and requests to Mission Support: Loft tank to be depleted to allow cleaning during crew changeover. Opportunity rover fully charged, however, the battery light is flashing and there is a bubbling sound coming from where the battery is located. It has been unplugged as a precaution.

Copyright © The Mars Society. All rights reserved. | Main Site